OpenStudy (anonymous):

14 Ariaroseb Write an equation of the line that passes through the given point and is parallel to the given line. 9. (-1,3); y=2x+2 10. (1,2)/ y= -5x+4 11. (18,2); 3y-x= -12 12. (2, -5); 2y=3x+10

OpenStudy (anonymous):

please help

OpenStudy (godlovesme):

OK use the point-slope formula which is y-y1=m(x-x1) M is ur slope while x1 is or given x coordinate and y1 is ur y coordinate Make sense?

OpenStudy (anonymous):

not really...

OpenStudy (godlovesme):

Srry =0 let's do 9 together we have y= 2x+2 so we have slope of 2 and y-intercept of 2 Y=MX+b r u familiar with this formula?

OpenStudy (anonymous):

yeah

OpenStudy (godlovesme):

Great! :) parallel lines have the same slope Y-y1=m(x-x1) we have (-1,3) -1 is x coordinate and 3 is y coordinate and we have slope of 2( we don't change the slope cuz parallels have same slope) So far so good?

OpenStudy (anonymous):

yeah I think so :)

OpenStudy (godlovesme):

Nice then let's use our point-slope formula y-y1=m(x-x1) M is ur slope while x1 is or given x coordinate and y1 is ur y coordinate This will give us Y-3= 2(x-(-1) Y-3= 2x +2 then adid 3 to bother side and tell me what u get :)

OpenStudy (anonymous):

Y-3=2x+5...? That's probably not right but I tried my best

OpenStudy (godlovesme):

U r doing it right but if u add 3 to both side it will be like this --> y-3+3=2x+2+3 -3+3 is 0 got it ?

OpenStudy (anonymous):

yeah

OpenStudy (godlovesme):

So our equation for #9 is y=2x+5 can u try that rest? Tag me if u get stuck :)

OpenStudy (anonymous):

Thank you

OpenStudy (godlovesme):

My pleasure! :D thanks for the medal

OpenStudy (godlovesme):

@mathmate & @ariaroseb thanks for the medal ^_^

OpenStudy (mathmate):

and a well-deserved one too! :)

OpenStudy (godlovesme):

Thanks:))